# Global well-posedness of the elastic-viscous-plastic sea-ice model with the inviscid Voigt-regularisation

### 06 Mar 2026 v2

@@ 3.2 Uniform-in-$(N,\epsilon)$ estimates
Hence, equations (50) and (51) yield
$\begin{gathered}\sup_{t\in[0,T]}\biggl[\lVert u_{N}(t)\rVert_{H^{2}}+\lVert\sigma_{N}(t)\rVert_{H^{3}}+\|\partial_{t}u_{N}(t)\|_{H^{2}}+\|\partial_{t}\sigma_{N}\|_{H^{3}}\biggr]\\ \leq C\exp(\exp(\exp(CT))),\end{gathered}$ (52)
@@
3.3 Well-posedness of the intermediate system (13)
Passing the limit $N\rightarrow\infty$ and existence of the strong solution. For any fixed $T\in(0,\infty)$ and $\epsilon\in(0,\infty)$, from the bound (52) and by applying the Aubin-Lions compactness lemma (see Theorem II.5.16 in Ref. [9]), one has the following convergence results: for a suitable subsequence of $N\rightarrow\infty$,
$u_{N}$ $\overset{\ast}{\rightharpoonup}u$ weakly-* in $~L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (53)
$\partial_{t}u_{N}$ $\overset{\ast}{\rightharpoonup}\partial_{t}u$ weakly-* in $~L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (54)
$u_{N}$ $\rightarrow u$ strongly in $~C([0,T];H^{1}(\mathbb{T}^{2})),$ (55)
$\sigma_{N}$ $\overset{\ast}{\rightharpoonup}\sigma$ weakly-* in $~L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (56)
$\sigma_{N}$ $\rightarrow\sigma$ strongly in $~C([0,T];H^{2}(\mathbb{T}^{2})),$ (57)
$\partial_{t}\sigma_{N}$ $\overset{\ast}{\rightharpoonup}\partial_{t}\sigma$ weakly-* in $~L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (58)
where $u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\partial_{t}u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$, and $\partial_{t}\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$ with the same estimates as in (52) with $(u_{N},\sigma_{N})$ replaced by $(u,\sigma)$, i.e.,
$\begin{gathered}\sup_{t\in[0,T]}\biggl[\lVert u(t)\rVert_{H^{2}}+\lVert\sigma(t)\rVert_{H^{3}}+\|\partial_{t}u(t)\|_{H^{2}}+\|\partial_{t}\sigma\|_{H^{3}}\biggr]\\ \leq C\exp(\exp(\exp(CT))),\end{gathered}$ (59)
for some constant $C$ independent of $\epsilon$ and $T$, but depending on $\mathcal{E},\alpha$, and the initial data. We also have $u\in C([0,T];H^{2}(\mathbb{T}^{2}))$ and $\sigma\in C([0,T];H^{3}(\mathbb{T}^{2}))$ thanks to Lemma 1.3.2 on p. 196 in Ref. [70].
It is clear that one can pass to the limit in the linear terms of the Galerkin system (25). To show the convergence of the nonlinear terms, first we observe that $\mathcal{D}_{\epsilon}^{N}\rightarrow\mathcal{D}_{\epsilon}$ in $C([0,T];L^{2}(\mathbb{T}^{2}))$. In fact, due to Lemma 2.3, we have by using (55)
@@ 3.3 Well-posedness of the intermediate system (13)
Notice that this bound holds pointwise in time, as will generally be the case for the estimates in this part of the paper. We will therefore omit the temporal part of the norm in this section. Next, by using (57) one can estimate that
@@
Therefore all the terms in the Galerkin system (25) converge, and $(u,\sigma)$ solves the intermediate system (13). This concludes the proof of the existence of a strong solution to the intermediate model (13).
@@
Suppose that $(u_{1},\sigma_{1}),(u_{2},\sigma_{2})\in C([0,T];H^{2}(\mathbb{T}^{2}))\times C([0,T];H^{3}(\mathbb{T}^{2}))$ are two strong solutions to the intermediate model (13). Let
@@
$\partial_{t}\delta u=\nabla\cdot\delta\sigma+\mathcal{T}_{w,1}-\mathcal{T}_{w,2}+\Omega(\delta u)^{\perp},$ (60a)
$\frac{1}{\mathcal{E}}\partial_{t}\big(\delta\sigma-\alpha^{2}\Delta\delta\sigma\big)+\frac{4\mathcal{D}_{\epsilon,1}}{P}(\sigma_{1}-\frac{1}{2}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2})-\frac{4\mathcal{D}_{\epsilon,2}}{P}(\sigma_{2}-\frac{1}{2}\operatorname{Tr}\sigma_{2}$ $\mathbb{I}_{2})$ (60b)
$+\frac{\mathcal{D}_{\epsilon,1}}{2P}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2}-\frac{\mathcal{D}_{\epsilon,2}}{2P}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2}+\frac{\mathcal{D}_{\epsilon,1}-\mathcal{D}_{\epsilon,2}}{2}$ $\mathbb{I}_{2}=D(\delta u),$ (60c)
@@
$\mathcal{T}_{w,i}$ $=c_{w}\rho_{w}\lvert U_{w}-u_{i}\rvert\big[(U_{w}-u_{i})\cos\theta+(U_{w}-u_{i})^{\perp}\sin\theta\big],$ (60d)
$\mathcal{D}_{\epsilon,i}$ $=\sqrt{\lvert D(u_{i})\rvert^{2}+\epsilon^{2}}.$ (60e)
We observe that $\partial_{t}(\delta u)\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$ and $\partial_{t}(\delta\sigma)\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$, and therefore we can take the $L^{2}$-inner product of equations (60a) and (60b) with $\delta u$ respectively $\delta\sigma$. We therefore compute the $L^{2}$-estimates on the difference $(\delta u,\delta\sigma)$ as follows (using a cancellation similar to (31))
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}\bigg]=-\int_{\mathbb{T}^{2}}\frac{\mathcal{D}_{\epsilon,1}-\mathcal{D}_{\epsilon,2}}{2}\operatorname{Tr}\delta\sigma\ {\rm d}{x}\\ -\int_{\mathbb{T}^{2}}\bigg[\frac{4\mathcal{D}_{\epsilon,1}}{P}(\sigma_{1}-\frac{1}{2}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2})-\frac{4\mathcal{D}_{\epsilon,2}}{P}(\sigma_{2}-\frac{1}{2}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2})\bigg]:\delta\sigma\ {\rm d}{x}\\ -\int_{\mathbb{T}^{2}}\bigg[\frac{\mathcal{D}_{\epsilon,1}}{2P}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2}-\frac{\mathcal{D}_{\epsilon,2}}{2P}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2}\bigg]:\delta\sigma\ {\rm d}{x}\\ +\int_{\mathbb{T}^{2}}\big(\mathcal{T}_{w,1}-\mathcal{T}_{w,2}\big)\cdot\delta u\ {\rm d}{x}=:J_{1}+J_{2}+J_{3}+J_{4}.\end{gathered}$ (61)
Applying Hölder’s inequality, inequality (21), and the Sobolev embedding theorem yields
@@
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}\bigg]\lesssim(\lVert\sigma_{1}\rVert_{L^{\infty}}+1)\big[\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\lVert\delta\sigma\rVert_{L^{2}}^{2}\big]\\ +\lVert\nabla u_{2}\rVert_{L^{2}}\big[\lVert\delta\sigma\rVert_{L^{2}}^{2}+\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}\big]+(\lVert u_{1}\rVert_{L^{\infty}}+\lVert u_{2}\rVert_{L\infty}+\lVert U_{w}\rVert_{L^{\infty}})\lVert\delta u\rVert_{L^{2}}^{2}.\end{gathered}$ (62)
Meanwhile, to calculate the $H^{1}$ estimate of the difference $(\delta u,\delta\sigma)$, similar to (42), we have (by using a cancellation similar to (31))
$\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg]$ (63)
@@
Similarly to the $L^{2}$-estimate, we establish (by using inequality (21) and the Sobolev embedding theorem)
@@
Finally, we estimate $J_{8}$. Recalling the notation (33), we denote by
$V_{w,i}\coloneqq U_{w}-u_{i},\qquad i=1,2,$ (64)
and hence $V_{w,2}-V_{w,1}=\delta u$. Thus, by using identity (35), one can compute, for $i=1,2$,
@@
$\|\partial_{i}(\mathcal{T}_{w,1}-\mathcal{T}_{w,2})\|_{L^{2}}\lesssim(\|V_{w,1}\|_{L^{\infty}}+\|V_{w,2}\|_{L^{\infty}})\|\partial_{i}\delta u\|_{L^{2}}+(\|\partial_{i}V_{w,1}\|_{L^{4}}+\|\partial_{i}V_{w,2}\|_{L^{4}})\|\delta u\|_{L^{4}}.$ (65)
Hence one has by using (64)
$|J_{8}|$ $\lesssim\lVert\nabla\delta u\rVert_{L^{2}}\bigg[\big(\lVert U_{w}-u_{1}\rVert_{L^{\infty}}+\lVert U_{w}-u_{2}\rVert_{L^{\infty}}\big)\lVert\nabla\delta u\rVert_{L^{2}}$ (66)
@@
Thus, we have derived the following from estimate (63),
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg]\\ \lesssim(1+\|\sigma_{1}\|_{H^{2}}+\|\sigma_{2}\|_{H^{2}}+\|u_{1}\|_{H^{2}}+\|u_{2}\|_{H^{2}})(\|\delta u\|_{H^{1}}^{2}+\|\delta\sigma\|_{H^{2}}^{2}).\end{gathered}$ (67)
Consequently, combining estimates (62) and (67) leads to
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\delta u\rVert_{L^{2}}^{2}+\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}(1+\alpha^{2})\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg]\\ \lesssim(1+\|\sigma_{1}\|_{H^{2}}+\|\sigma_{2}\|_{H^{2}}+\|u_{1}\|_{H^{2}}+\|u_{2}\|_{H^{2}})(\|\delta u\|_{H^{1}}^{2}+\|\delta\sigma\|_{H^{2}}^{2}),\end{gathered}$ (68)
from which the stability (i.e. the continuous dependence on the initial data) and uniqueness of the strong solution follows. This concludes the proof of the global well-posedness of the intermediate model (13).
@@
We consider a sequence of solutions $\{(u_{\epsilon},\sigma_{\epsilon})\}$ of the intermediate system (13) such that $\epsilon\rightarrow 0$. Thanks to estimate (59), which is independent of $\epsilon$, by the Aubin-Lions lemma and Banach-Alaoglu theorem we have the following convergence results
$u_{\epsilon}$ $\overset{\ast}{\rightharpoonup}u$ weakly-* in $~L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (69)
$\partial_{t}u_{\epsilon}$ $\overset{\ast}{\rightharpoonup}\partial_{t}u$ weakly-* in $~L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (70)
$u_{\epsilon}$ $\rightarrow u$ strongly in $~C([0,T];H^{1}(\mathbb{T}^{2})),$ (71)
$\sigma_{\epsilon}$ $\overset{\ast}{\rightharpoonup}\sigma$ weakly-* in $~L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (72)
$\sigma_{\epsilon}$ $\rightarrow\sigma$ strongly in $~C([0,T];H^{2}(\mathbb{T}^{2})),$ (73)
$\partial_{t}\sigma_{\epsilon}$ $\overset{\ast}{\rightharpoonup}\partial_{t}\sigma$ weakly-* in $~L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (74)
where the limits $u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\partial_{t}u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$, and $\partial_{t}\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$ satisfy estimate (59). It is left to verify that $(u,\sigma)$ is a solution to the Voigt-EVP model (12).
@@
where we have used Lemma 2.3, the convergence $u_{\epsilon}\rightarrow u$ in $C([0,T];H^{1}(\mathbb{T}^{2}))$ and the elementary inequality $\sqrt{y+z}\leq\sqrt{y}+\sqrt{z}$ for $y,z\geq 0$. It is clear that the linear terms in equations (13a)-(13b) converge. We now show that the nonlinear terms in the equation for the stress tensor converge. In particular, for all $t\in[0,T]$ we have
@@
$\partial_{t}\delta u=\nabla\cdot\delta\sigma+\mathcal{T}_{w,1}-\mathcal{T}_{w,2}+\Omega(\delta u)^{\perp},$ (75a)
$\frac{1}{\mathcal{E}}\partial_{t}\big(\delta\sigma-\alpha^{2}\Delta\delta\sigma\big)+\frac{4\mathcal{D}_{1}}{P}(\sigma_{1}-\frac{1}{2}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2})-\frac{4\mathcal{D}_{2}}{P}(\sigma_{2}-\frac{1}{2}\operatorname{Tr}\sigma_{2}$ $\mathbb{I}_{2})$ (75b)
$+\frac{\mathcal{D}_{1}}{2P}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2}-\frac{\mathcal{D}_{2}}{2P}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2}+\frac{\mathcal{D}_{1}-\mathcal{D}_{2}}{2}$ $\mathbb{I}_{2}=D(\delta u),$ (75c)

### 20 Mar 2026 v3

@@ 3.2 Uniform-in-$(N,\epsilon)$ estimates
Hence, equations (3.26) and (3.27) yield
$\begin{gathered}\sup_{t\in[0,T]}\biggl[\lVert u_{N}(t)\rVert_{H^{2}}+\lVert\sigma_{N}(t)\rVert_{H^{3}}+\|\partial_{t}u_{N}(t)\|_{H^{2}}+\|\partial_{t}\sigma_{N}\|_{H^{3}}\biggr]\\ \leq C\exp(\exp(\exp(CT))),\end{gathered}$ (3.28)
@@
3.3 Well-posedness of the intermediate system (1.13)
Passing the limit $N\rightarrow\infty$ and existence of the strong solution. For any fixed $T\in(0,\infty)$ and $\epsilon\in(0,\infty)$, from the bound (3.28) and by applying the Aubin-Lions compactness lemma (see [9, Theorem II.5.16]), one has the following convergence results: for a suitable subsequence of $N\rightarrow\infty$,
$u_{N}$ $\overset{\ast}{\rightharpoonup}u$ weakly-* in $~L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (3.29)
$\partial_{t}u_{N}$ $\overset{\ast}{\rightharpoonup}\partial_{t}u$ weakly-* in $~L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (3.30)
$u_{N}$ $\rightarrow u$ strongly in $~C([0,T];H^{1}(\mathbb{T}^{2})),$ (3.31)
$\sigma_{N}$ $\overset{\ast}{\rightharpoonup}\sigma$ weakly-* in $~L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (3.32)
$\sigma_{N}$ $\rightarrow\sigma$ strongly in $~C([0,T];H^{2}(\mathbb{T}^{2})),$ (3.33)
$\partial_{t}\sigma_{N}$ $\overset{\ast}{\rightharpoonup}\partial_{t}\sigma$ weakly-* in $~L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (3.34)
where $u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\partial_{t}u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$, and $\partial_{t}\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$ with the same estimates as in (3.28) with $(u_{N},\sigma_{N})$ replaced by $(u,\sigma)$, i.e.,
$\begin{gathered}\sup_{t\in[0,T]}\biggl[\lVert u(t)\rVert_{H^{2}}+\lVert\sigma(t)\rVert_{H^{3}}+\|\partial_{t}u(t)\|_{H^{2}}+\|\partial_{t}\sigma\|_{H^{3}}\biggr]\\ \leq C\exp(\exp(\exp(CT))),\end{gathered}$ (3.35)
for some constant $C$ independent of $\epsilon$ and $T$, but depending on $\mathcal{E},\alpha$, and the initial data. We also have $u\in C([0,T];H^{2}(\mathbb{T}^{2}))$ and $\sigma\in C([0,T];H^{3}(\mathbb{T}^{2}))$ thanks to [70, Lemma 1.3.2, p. 196].
It is clear that one can pass to the limit in the linear terms of the Galerkin system (3.1). To show the convergence of the nonlinear terms, first we observe that $\mathcal{D}_{\epsilon}^{N}\rightarrow\mathcal{D}_{\epsilon}$ in $C([0,T];L^{2}(\mathbb{T}^{2}))$. In fact, due to Lemma 2.3, we have by using (3.31)
@@
Notice that this bound holds pointwise in time, as will generally be the case for the estimates in this part of the paper. We will therefore omit the temporal part of the norm in this section. Next, by using (3.33) one can estimate that
@@
Therefore all the terms in the Galerkin system (3.1) converge, and $(u,\sigma)$ solves the intermediate system (1.13). This concludes the proof of the existence of a strong solution to the intermediate model (1.13).
@@
Suppose that $(u_{1},\sigma_{1}),(u_{2},\sigma_{2})\in C([0,T];H^{2}(\mathbb{T}^{2}))\times C([0,T];H^{3}(\mathbb{T}^{2}))$ are two strong solutions to the intermediate model (1.13). Let
@@
$\partial_{t}\delta u=\nabla\cdot\delta\sigma+\mathcal{T}_{w,1}-\mathcal{T}_{w,2}+\Omega(\delta u)^{\perp},$ (3.36a)
$\frac{1}{\mathcal{E}}\partial_{t}\big(\delta\sigma-\alpha^{2}\Delta\delta\sigma\big)+\frac{4\mathcal{D}_{\epsilon,1}}{P}(\sigma_{1}-\frac{1}{2}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2})-\frac{4\mathcal{D}_{\epsilon,2}}{P}(\sigma_{2}-\frac{1}{2}\operatorname{Tr}\sigma_{2}$ $\mathbb{I}_{2})$ (3.36b)
$+\frac{\mathcal{D}_{\epsilon,1}}{2P}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2}-\frac{\mathcal{D}_{\epsilon,2}}{2P}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2}+\frac{\mathcal{D}_{\epsilon,1}-\mathcal{D}_{\epsilon,2}}{2}$ $\mathbb{I}_{2}=D(\delta u),$ (3.36c)
@@
$\mathcal{T}_{w,i}$ $=c_{w}\rho_{w}\lvert U_{w}-u_{i}\rvert\big[(U_{w}-u_{i})\cos\theta+(U_{w}-u_{i})^{\perp}\sin\theta\big],$ (3.36d)
$\mathcal{D}_{\epsilon,i}$ $=\sqrt{\lvert D(u_{i})\rvert^{2}+\epsilon^{2}}.$ (3.36e)
We observe that $\partial_{t}(\delta u)\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$ and $\partial_{t}(\delta\sigma)\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$, and therefore we can take the $L^{2}$-inner product of equations (3.36a) and (3.36b) with $\delta u$ respectively $\delta\sigma$. We therefore compute the $L^{2}$-estimates on the difference $(\delta u,\delta\sigma)$ as follows (using a cancellation similar to (3.7))
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}\bigg]=-\int_{\mathbb{T}^{2}}\frac{\mathcal{D}_{\epsilon,1}-\mathcal{D}_{\epsilon,2}}{2}\operatorname{Tr}\delta\sigma\ {\rm d}{x}\\ -\int_{\mathbb{T}^{2}}\bigg[\frac{4\mathcal{D}_{\epsilon,1}}{P}(\sigma_{1}-\frac{1}{2}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2})-\frac{4\mathcal{D}_{\epsilon,2}}{P}(\sigma_{2}-\frac{1}{2}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2})\bigg]:\delta\sigma\ {\rm d}{x}\\ -\int_{\mathbb{T}^{2}}\bigg[\frac{\mathcal{D}_{\epsilon,1}}{2P}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2}-\frac{\mathcal{D}_{\epsilon,2}}{2P}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2}\bigg]:\delta\sigma\ {\rm d}{x}\\ +\int_{\mathbb{T}^{2}}\big(\mathcal{T}_{w,1}-\mathcal{T}_{w,2}\big)\cdot\delta u\ {\rm d}{x}=:J_{1}+J_{2}+J_{3}+J_{4}.\end{gathered}$ (3.37)
Applying Hölder’s inequality, inequality (2.7), and the Sobolev embedding theorem yields
@@
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}\bigg]\lesssim(\lVert\sigma_{1}\rVert_{L^{\infty}}+1)\big[\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\lVert\delta\sigma\rVert_{L^{2}}^{2}\big]\\ +\lVert\nabla u_{2}\rVert_{L^{2}}\big[\lVert\delta\sigma\rVert_{L^{2}}^{2}+\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}\big]+(\lVert u_{1}\rVert_{L^{\infty}}+\lVert u_{2}\rVert_{L\infty}+\lVert U_{w}\rVert_{L^{\infty}})\lVert\delta u\rVert_{L^{2}}^{2}.\end{gathered}$ (3.38)
Meanwhile, to calculate the $H^{1}$ estimate of the difference $(\delta u,\delta\sigma)$, similar to (3.18), we have (by using a cancellation similar to (3.7))
$\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg]$ (3.39)
@@
Similarly to the $L^{2}$-estimate, we establish (by using inequality (2.7) and the Sobolev embedding theorem)
@@
Finally, we estimate $J_{8}$. Recalling the notation (3.9), we denote by
$V_{w,i}\coloneqq U_{w}-u_{i},\qquad i=1,2,$ (3.40)
and hence $V_{w,2}-V_{w,1}=\delta u$. Thus, by using identity (3.11), one can compute, for $i=1,2$,
@@
$\|\partial_{i}(\mathcal{T}_{w,1}-\mathcal{T}_{w,2})\|_{L^{2}}\lesssim(\|V_{w,1}\|_{L^{\infty}}+\|V_{w,2}\|_{L^{\infty}})\|\partial_{i}\delta u\|_{L^{2}}+(\|\partial_{i}V_{w,1}\|_{L^{4}}+\|\partial_{i}V_{w,2}\|_{L^{4}})\|\delta u\|_{L^{4}}.$ (3.41)
Hence one has by using (3.40)
$|J_{8}|$ $\lesssim\lVert\nabla\delta u\rVert_{L^{2}}\bigg[\big(\lVert U_{w}-u_{1}\rVert_{L^{\infty}}+\lVert U_{w}-u_{2}\rVert_{L^{\infty}}\big)\lVert\nabla\delta u\rVert_{L^{2}}$ (3.42)
@@
Thus, we have derived the following from estimate (3.39),
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg]\\ \lesssim(1+\|\sigma_{1}\|_{H^{2}}+\|\sigma_{2}\|_{H^{2}}+\|u_{1}\|_{H^{2}}+\|u_{2}\|_{H^{2}})(\|\delta u\|_{H^{1}}^{2}+\|\delta\sigma\|_{H^{2}}^{2}).\end{gathered}$ (3.43)
Consequently, combining estimates (3.38) and (3.43) leads to
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\delta u\rVert_{L^{2}}^{2}+\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}(1+\alpha^{2})\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg]\\ \lesssim(1+\|\sigma_{1}\|_{H^{2}}+\|\sigma_{2}\|_{H^{2}}+\|u_{1}\|_{H^{2}}+\|u_{2}\|_{H^{2}})(\|\delta u\|_{H^{1}}^{2}+\|\delta\sigma\|_{H^{2}}^{2}),\end{gathered}$ (3.44)
from which the stability (i.e. the continuous dependence on the initial data) and uniqueness of the strong solution follows. This concludes the proof of the global well-posedness of the intermediate model (1.13).
@@
We consider a sequence of solutions $\{(u_{\epsilon},\sigma_{\epsilon})\}$ of the intermediate system (1.13) such that $\epsilon\rightarrow 0$. Thanks to estimate (3.35), which is independent of $\epsilon$, by the Aubin-Lions lemma and Banach-Alaoglu theorem we have the following convergence results
$u_{\epsilon}$ $\overset{\ast}{\rightharpoonup}u$ weakly-* in $~L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (4.1)
$\partial_{t}u_{\epsilon}$ $\overset{\ast}{\rightharpoonup}\partial_{t}u$ weakly-* in $~L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (4.2)
$u_{\epsilon}$ $\rightarrow u$ strongly in $~C([0,T];H^{1}(\mathbb{T}^{2})),$ (4.3)
$\sigma_{\epsilon}$ $\overset{\ast}{\rightharpoonup}\sigma$ weakly-* in $~L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (4.4)
$\sigma_{\epsilon}$ $\rightarrow\sigma$ strongly in $~C([0,T];H^{2}(\mathbb{T}^{2})),$ (4.5)
$\partial_{t}\sigma_{\epsilon}$ $\overset{\ast}{\rightharpoonup}\partial_{t}\sigma$ weakly-* in $~L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (4.6)
where the limits $u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\partial_{t}u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$, and $\partial_{t}\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$ satisfy estimate (3.35). It is left to verify that $(u,\sigma)$ is a solution to the Voigt-EVP model (1.12).
@@
where we have used Lemma 2.3, the convergence $u_{\epsilon}\rightarrow u$ in $C([0,T];H^{1}(\mathbb{T}^{2}))$ and the elementary inequality $\sqrt{y+z}\leq\sqrt{y}+\sqrt{z}$ for $y,z\geq 0$. It is clear that the linear terms in equations (1.13a)-(1.13b) converge. We now show that the nonlinear terms in the equation for the stress tensor converge. In particular, for all $t\in[0,T]$ we have
@@
$\partial_{t}\delta u=\nabla\cdot\delta\sigma+\mathcal{T}_{w,1}-\mathcal{T}_{w,2}+\Omega(\delta u)^{\perp},$ (4.7a)
$\frac{1}{\mathcal{E}}\partial_{t}\big(\delta\sigma-\alpha^{2}\Delta\delta\sigma\big)+\frac{4\mathcal{D}_{1}}{P}(\sigma_{1}-\frac{1}{2}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2})-\frac{4\mathcal{D}_{2}}{P}(\sigma_{2}-\frac{1}{2}\operatorname{Tr}\sigma_{2}$ $\mathbb{I}_{2})$ (4.7b)
$+\frac{\mathcal{D}_{1}}{2P}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2}-\frac{\mathcal{D}_{2}}{2P}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2}+\frac{\mathcal{D}_{1}-\mathcal{D}_{2}}{2}$ $\mathbb{I}_{2}=D(\delta u),$ (4.7c)